OpenStudy (anonymous):

An acute angle θ is in a right triangle with cos θ = 9/10 . What is the value of sec θ?

OpenStudy (ghazi):

sec theta= 1/cos theta

OpenStudy (anonymous):

9/19

OpenStudy (ghazi):

nope....it would be 10/9

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

ohh wait nvm i sees

OpenStudy (ghazi):

sec theta= 1/cos theta....now substitute the value of cos theta you'll get 1/9/10= 10/9